Each firm in a perfectly competitive market has long run average cost represented as AC(q) =...

  1. Each firm in a perfectly competitive market has long run average cost represented as AC(q) = 100q- 10+100/q. Long run marginal cost is MC=200q-10. The market demand is Qd = 2150-5P.    Find the long run equilibrium output per firm, q*, the long run equilibrium price, P*, and the number of firms in the industry, n*.

P = 190; Q = 1200; q =1 , n = 1200
Answer #1

In long run equilibrium, AC = MC = P.

100q - 10 + (100/q) = 200q - 10

100q = 100/q

q2 = 1

q = 1

P = MC = (200 x 1) - 10 = 200 - 10 = 190

Qd = 2150 - (5 x 190) = 2150 - 950 = 1200

n = Qd / q = 1200/1 = 1200